(i) If 4𝑒 π‘₯ βˆ’ 3𝑒 βˆ’π‘₯ ≑ π‘ƒπ‘ π‘–π‘›β„Ž π‘₯ + π‘„π‘π‘œπ‘ β„Ž π‘₯, determine the values of 𝑃 and 𝑄. (ii) Solve the equation π‘π‘œπ‘ π‘’π‘β„Ž 𝑦 = βˆ’0.4458, correct to 4 significant figures


1
Expert's answer
2021-09-29T00:01:34-0400

We have

"sinh(x)=\frac{e^x-e^{-x}}{2}\\
cosh(x)=\frac{e^x-+e^{-x}}{2}\\
P\cdot sinh(x)+Q\cdot xosh(x)=P\cdot \left( \frac{e^x-e^{-x}}{2} \right)+
Q\cdot \left( \frac{e^+-e^{-x}}{2} \right)=\\
e^x\cdot \left( \frac{P+Q}{2} \right)+
e^{-x}\cdot \left( \frac{-P+Q}{2} \right)=4\cdot e^x-3\cdot e^{-x}"

Therefore

"\begin{cases}
 P+Q=8 \\
 -P+Q=-6
\end{cases}"

Adding equatioms we have 2Q=2=>Q=1.

Inserting this value to the first equation have P=8-1=7;

Answer: P=7,Q=1

2)cosech(x)=1/sinh(x)=

"\frac{2}{e^x-e^{-x}}=-0.4458;\\
e^x-e^{-x}=\frac{2}{-0.4458}=-4.486\\"

Let be "e^x=t"

Then we have an equation

"t-\frac{1}{t}=-4.486\\
t^2+4.486\cdot t-1=0;\\
t=\frac{-4.486+\sqrt{4.486^2+4}}{2}=0.2128"

"e^x=0.2128"

x=ln(0.2128)=-1.5474

Answer: x=-1.5474 (or y=-1.5474)